OpenStudy (anonymous):

f(x)=radical(8-e^(2x)). Find domain of f? inverse of f? and domain of inverse of f?

OpenStudy (anonymous):

\[f(x)=\sqrt{8-e^{2x}}\] you have to be sure that \[8-e^{2x}\geq 0\] meaning \[e^{2x}\leq 8\] \[2x\leq \ln(8)\] \[x\leq \frac{\ln(8)}{2}\]

OpenStudy (anonymous):

if you want to be fancy you can write it as \[x\leq \frac{3}{2}\ln(2)\]

OpenStudy (anonymous):

to find the inverse maybe write \[x=\sqrt{8-e^{2y}}\] and solve for y via \[x^2=8-e^{2y}\] \[x^2-8=-e^{2y}\] \[e^{2y}=8-x^2\] \[2y=\ln(8-x^2)\] \[y=\frac{1}{2}\ln(8-x^2)\] \[y=\ln(\sqrt{8-x^2})\]

OpenStudy (anonymous):

what about domain? and to enter it there are two boxes with a comma in between

OpenStudy (anonymous):

you have to make sure that \[8-x^2>0\] because you cannot take the square root of a negative number and you cannot take the log of 0

OpenStudy (anonymous):

oh hold on now i have to think. the range of f is non - negative, so you cannot use a negative number for the inverse

myininaya (myininaya):

satellite what are you thinking about?

OpenStudy (anonymous):

so domain is \[[0,\sqrt{8})\] or \[[0,2\sqrt{2})\]

OpenStudy (anonymous):

oh i just had to make sure the answer made sense

OpenStudy (anonymous):

if you just see \[\frac{1}{2}\ln(8-x^2)\] you would say domain is \[(-\sqrt{8},\sqrt{8})\] but not in this case

myininaya (myininaya):

\[f(x)=\sqrt{8-e^{2x}}\] domain of f is \[8-e^{2x} \ge 0 \] \[8 \ge e^{2x}\] natural log of both sides \[\ln(8) \ge \ln(e^{2x})\] \[\ln(8) \ge 2x \ln(e)\] \[\ln(8) \ge 2x\] \[\frac{\ln(8)}{2} \ge x\]

OpenStudy (anonymous):

yeah that is the first thing i wrote

OpenStudy (anonymous):

then i found the inverse. but i had to think for a moment about the domain of the inverse

myininaya (myininaya):

\[y=\sqrt{8-e^{2x}}\] square both sides \[y^2=8-e^{2x}\] subtract 8 on both sides \[y^2-8=-e^{2x}\] multiply negative one on both sides \[-y^2+8=e^{2x}\] take natural log of both sides \[\ln(-y^2+8)=\ln(e^{2x})\] \[\ln(8-y^2)=2x \ln(e)\] \[\ln(8-y^2)=2x\] divide both sides by 2 \[\frac{\ln(8-y^2)}{2}=x\] \[f^{-1}(x)=\frac{\ln(8-x^2)}{2}\]

OpenStudy (anonymous):

yeah i did that too!

myininaya (myininaya):

\[8-x^2>0\] is the domain

OpenStudy (anonymous):

that is where the mistake is!

myininaya (myininaya):

why?

OpenStudy (anonymous):

i wrote it out up top. what you wrote i mean. but the domain of \[f^{-1}(x)=\frac{1}{2}\ln(8-x^2)=\ln(\sqrt{8-x^2})\] is in fact \[[0,\sqrt{8})\]

OpenStudy (anonymous):

because it is defined as the inverse of \[f(x)=\sqrt{8-e^{2x}}\] and the range of this function does not include negative numbers. so the domain of the inverse does not that is why i had to take a moment to think

myininaya (myininaya):

yes you are right! :)

OpenStudy (anonymous):

imagine!

myininaya (myininaya):

so we would say the domain of f inverse is 8-x^2>0 with also remembering that we must also satisfy x>=0

OpenStudy (anonymous):

remember... \cup also \cap \[\color{red}{\text{red}}\cap \color{blue}{\text{blue}}\]

OpenStudy (anonymous):

yeah domain is \[[0,2\sqrt{2})\]

myininaya (myininaya):

first we find that \[(-\sqrt{8},\sqrt{8})\] from 8-x^2>0 but then we remember that we must have x>=0 so we only actually have \[[0,\sqrt{8})\]

myininaya (myininaya):

and yes \[\sqrt{8}=\sqrt{4 \cdot 2}=2\sqrt{2}\]

OpenStudy (anonymous):

okay but what about the domain of f? admist all this i got lost a little

myininaya (myininaya):

the domain of f is the first thing i found

OpenStudy (anonymous):

ln(8)/2? how do i write that in two boxes with a comma in between

myininaya (myininaya):

are you saying how you write an interval notation? (-inf,ln(8)/2]